# Ferromagnetism Mechanism in a Geometrically Frustrated Triangular Lattice

### v1

@@ I DMRG Details and DMRG Convergence
Fig. S1: Line-cut plot of the static spin structure $S(\mathbf{q})$ along the momentum path passing through the $\Gamma$ point (denoted as the white dashed line in Fig. 1(c) in the main text) with respect to various bond dimensions $D$.
In this section, we provide more numerical details in our calculations, as well as DMRG convergence with bond dimension. We study the ground-state properties of the triangular-lattice Hubbard model using DMRG. The Hamiltonian reads
$H=-t\sum_{\langle\mathbf{i}\mathbf{j}\rangle,\sigma}\left(c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}+\text{H.c.}\right)+U\sum_{\mathbf{i}}n_{\mathbf{i},\uparrow}n_{\mathbf{i},\downarrow}.$ (S1)
[…]
We employ DMRG to simulate the Hamiltonian in Eq. (S1) on triangular lattice cylinders with primitive vectors $\bm{e}_{x}=(1,0)$, $\bm{e}_{y}=(1/2,\sqrt{3}/2)$, wrapping on cylinders with a lattice spacing of unity. System size is $N=L_{x}\times L_{y}$, where $L_{x}$ ($L_{y}$) represents the length of the cylinder (circumference). Here we mainly study cylinders with width $L_{y}=4$, $6$ and set the energy unit as $t$. In particular, we extend our computation from finite-$U$ to the infinite-$U$ limit using the effective $t$-$J$ model with vanishing superexchange interaction $J=4t^{2}/U$. Considering the varying convergence rates at different parameters, we set the DMRG bond dimension $D$ up to $D\approx 36,000$.
We provide specific examples to illustrate the necessity of large bond dimensions for accurately identifying the ground-state properties of both the ferromagnetic phase and incommensurate SDW phases. Additionally, we have performed DMRG calculations with different initial states to ensure the convergence of DMRG.
For incommensurate SDW region, at relatively small bond dimensions, the peak of the static spin structure $S(\mathbf{q})$ is located at the momentum $\mathbf{\Gamma}$ [see Fig. S1(a,c,d)], which is a signature for ferromagnetism. However, with the increase of bond dimension, the energy becomes lower and the peak splits around the momentum $\mathbf{\Gamma}$. These observations indicate large bond dimensions are required to obtain converged results for distinguishing the ferromagnetism and iSDW with close energies.
@@
On the other side, for the ferromagnetism region, a large bond dimension is crucial for determining whether the ground state is partially polarized or fully polarized. As presented in Fig. S1(b), the peak of $S(\mathbf{q})$ gets higher with increasing bond dimension, until it saturates. Notably, DMRG convergence becomes more challenging with decreasing interaction strength $U/t$ and increasing system size $N$. In Fig. S1(f,e), we also present the good convergence of the DMRG simulations in the iSDW region at $U=\infty$ that separates the ferromagnetic phases at $\delta\approx 1/16$ and $\delta\approx 1/2$ [see Fig. S6(b)].
II Details on unrestricted Hartree-Fock method
We consider the Hubbard model on the triangular lattice with the Hamiltonian
$H=-\sum_{\langle\mathbf{ij}\rangle,\sigma}(t+\mu\delta_{ij})c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}+U\sum_{i}n_{\mathbf{i}\uparrow}n_{\mathbf{i}\downarrow}$ (S2)
Here $t$ is the hopping integral and $U$ is the on-site Hubbard interaction strength. Our focus is on a scenario of high electron doping $\delta=0.5$ and applying the unrestricted Hartree-Fock (UHF) approximation. In the UHF approximation, the mean-field order parameters are introduced for both the on-site densities $\langle n_{\mathbf{i},\sigma}\rangle$ and the spin-flips $\langle S_{\mathbf{i}}^{-}\rangle$ and $\langle S_{\mathbf{i}}^{+}\rangle$. Then, the Hartree-Fock factorizes the Hamiltonian to
$H_{\mathrm{HF}}=$ $-\sum_{\langle\mathbf{ij}\rangle}(t+\mu\delta_{\mathbf{ij}})c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}+\sum_{i}U[\langle n_{\mathbf{i},\uparrow}\rangle n_{\mathbf{i},\downarrow}+\langle n_{\mathbf{i},\downarrow}\rangle n_{\mathbf{i},\uparrow}-\langle S_{\mathbf{i}}^{+}\rangle S_{\mathbf{i}}^{-}-\langle S_{\mathbf{i}}^{-}\rangle S_{\mathbf{i}}^{+}]+\sum_{\mathbf{i}}U\left[-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle\right],$
Here the mean-field parameters is an expectation value under the ground state $|\mathrm{HF}\rangle$ of the $H_{\mathrm{HF}}$ mean-field Hamiltonian,
$\langle n_{\mathbf{i},\sigma}\rangle\equiv\langle\mathrm{HF}|n_{\mathbf{i},\sigma}|\mathrm{HF}\rangle,\quad\langle S_{\mathbf{i}}^{\pm}\rangle\equiv\langle\mathrm{HF}|S_{\mathbf{i}}^{\pm}|\mathrm{HF}\rangle,$ (S3)
@@ II Details on unrestricted Hartree-Fock method
and the on-site densities are subject to the particle number constraint $\sum_{\mathbf{i}}[\langle n_{\mathbf{i},\uparrow}\rangle+\langle n_{\mathbf{i},\downarrow}\rangle]=N(1+\delta)$, with $N$ being the number of the lattice sites. We have total $4N$ variational parameters, which can be iteratively solved by the self-consistent equations
$\langle n_{\mathbf{i},\sigma}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|n_{\mathbf{i},\sigma}|\psi_{n}\rangle$ (S4)
$\langle S_{\mathbf{i}}^{\pm}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|S_{\mathbf{i}}^{\pm}|\psi_{n}\rangle$ (S5)
@@
Fig. S2: The static spin structure factor $S(\mathbf{q})$ as a function of $U/t$ at electron doping $\delta=1/2$. (a)-(c) The contour plot of $S(\mathbf{q})$ for the $N=24\times 6$ system at $U/t=4$ (a), $U/t=10$ (b), and $U/t=24$ (c). The black dots represent the accessible momenta in the Brillouin zone (denoted as a white hexagon), and interpolation has been applied in the contour plot. (d) Schematic of the evolution of the $\mathbf{q}_{0}$ with increasing $U/t$, indicated by the transition from lighter to darker colors following the black dashed arrows. Here, $\mathbf{q}_{0}$ denotes the momenta where $S(\mathbf{q})$ reaches its maximum, within the Brillouin zone represented by a hexagon. (e) Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ as a function of $U/t$. For larger $U/t$, the peak of $S(\mathbf{q})$ stabilizes at momentum $\mathbf{\Gamma}$ with a saturation value independent of $N$, indicating the fully polarized ferromagnetism. (f) Line-cut plot of $S(\mathbf{q})$ along the momentum path passing through the $\mathbf{\Gamma}$ point, depicted as a white dashed line in (c). (g) Polarization $P\equiv 2Nm_{s}(\mathbf{\Gamma})$ from the unrestricted Hartree-Fock method on $N=48\times 48$ and $N=24\times 24$ systems with periodic boundary conditions.
@@
For a stable convergence, we utilize the direct inversion in the iterative subspace (DIIS) method. In this approach, we can express the Hartree-Fock Hamiltonian in matrix form as follows:
@@
$H_{\mathrm{HF}}=C^{\dagger}FC+\sum_{\mathbf{i}}U\left[-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle\right]$ (S6)
where $C$ is a vector with $C_{\mathbf{i}\sigma}=c_{\mathbf{i},\sigma}$ and $F$ is the Fock matrix. For the magnetic order, the single-particle correlation matrix $\Phi$ is defined as
$\Phi_{\mathbf{ij},\alpha\beta}=\langle c_{\mathbf{i},\alpha}^{\dagger}c_{\mathbf{j},\beta}\rangle.$ (S7)
The DIIS method relies on the property that a convergence condition is met when the single-particle correlation matrix commutes with the Fock matrix
$F\Phi-\Phi F=0.$ (S8)
This condition allows us to define an error vector for the $n$-th iteration as
$\mathbf{e}_{n}=F_{n}\Phi_{n}-\Phi_{n}F_{n},$ (S9)
where $\Phi_{n}$ is obtained from the diagonalization of $F_{n}$. The iteration process involves updating the Fock matrix as
$F_{k}=\sum_{n=1}^{k-1}d_{n}F_{n}.$ (S10)
Here, $d_{n}$ are the DIIS coefficients obtained through a least-squares constrained minimization of the error vectors. The aim of the optimization is to minimize the error given by
$\mathrm{Error}$ $=|\sum_{n}d_{n}\mathbf{e}_{n}|^{2}$ (S11)
with the constraint $\sum_{n}d_{n}=1$. In our calculations, we set the convergent condition of the energy difference between two subsequent iteration processes as $10^{-8}t$.
III Itinerant Ferromagnetism at Electron Doping $\delta=1/2$
@@
Fig. S3: (a-d) The contour plot of the static spin structure factor $S(\mathbf{q})$ across a range of $U/t$ with $N=8\times 4$ at $\delta=1/2$ for the Hubbard model.
@@
Fig. S4: (a-d) The contour plot of the static spin structure factor $S(\mathbf{q})$ across a range of $U/t$ with $N=12\times 4$ at $\delta=1/2$ for the Hubbard model.
In the main text, we have presented key findings at electron doping $\delta=1/2$. In this section, we provide additional numerical details and numerical data. For the convenience of readers, we also show the data presented in the main text to gain a complete picture of the evolution of magnetism when tuning the coupling strength $U/t$ from finite to infinite $U/t$, as depicted in Fig. S2. These analyses reveal the emergence of itinerant ferromagnetism with increasing Coulomb interaction and enable a comprehensive understanding of the underlying mechanism under general interaction strength $U/t$.
The magnetic order is signaled by peaks at certain wave vectors $\mathbf{q}=\mathbf{q}_{0}$ in the static spin structure factor $S(\mathbf{q})$, defined by $S(\mathbf{q})\equiv\sum_{\mathbf{i},\mathbf{j}}\langle{\mathbf{S}_{\mathbf{i}}\cdot\mathbf{S}_{\mathbf{j}}}\rangle e^{i\mathbf{q}\cdot(\mathbf{i}-\mathbf{j})}/N$. We first focus on the evolution of the momenta $\mathbf{q}_{0}$ as a function of interaction strength $U/t$. For the noninteracting case ($U/t=0$) in a triangular lattice, electron doping $\delta=1/2$ corresponds to the Van-Hove filling at which a perfect nesting of the Fermi surface exists. In the weak-coupling limit $U/t\to 0$, it is expected that nesting leads to peaks of $S(\mathbf{q})$ locating at the $\mathbf{M}$ points in the Brillouin zone. Previous weak-coupling theory proposes that the candidate ground state exhibits noncoplanar chiral magnetic order. When the weak coupling assumption is no longer applicable at intermediate to strong interaction strengths $U/t$, using DMRG we numerically observe that other magnetic ordering vectors distinct from the nesting vectors emerge, as shown in Figs. S2(a), (b) for $L_{y}=6$. Interestingly, the locations of these peaks in $S(\mathbf{q})$ shift towards the $\mathbf{\Gamma}$ point as $U/t$ increases. Upon surpassing a critical strength $U_{c}/t$ toward the limit $U/t=\infty$, the peak of $S(\mathbf{q})$ becomes fixed at $\mathbf{\Gamma}$ point, as illustrated in Fig. S2(c), (e), indicating the establishment of a stabilized ferromagnetic phase in the strong coupling regime. In Fig. S2(d), we schematically plot such evolution of the peaks in $S(\mathbf{q})$ with increasing $U/t$, as visualized by the dashed arrow from momenta $\mathbf{M}$ to $\mathbf{\Gamma}$. Notably, after the peak of $S(\mathbf{q})$ shifts to the $\mathbf{\Gamma}$ point, its height gets larger in a finite range of increasing $U/t$ [see Fig. S2(f)], until reaching its maximum value and forming a plateau, as illustrated by Fig. S2(e). This signals the evolution from a partially-polarized to a fully-polarized phase. In particular, the saturated value matches the ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})\equiv S(\mathbf{\Gamma})/N\sim\delta_{\mathbf{q},\mathbf{\Gamma}}/16$ for a fully polarized state at electron doping $\delta=1/2$.
@@
We present additional supporting data for the evolution of the momentum $\mathbf{q}_{0}$ of the peak in $S(\mathbf{q})$ as a function of interaction strength $U/t$ in Fig. S3 for $N=8\times 4$ and Fig. S4 for $N=12\times 4$. These results for $L_{y}=4$ systems are consistent with the case of $L_{y}=6$ in Fig. S2(a)-(c). The peak locations in $S(\mathbf{q})$ progressively shift toward the $\mathbf{\Gamma}$ point as $U/t$ increases, as presented in Fig.S3(a-c) and Fig.S4(a,b). This observation supports the schematic plot in Fig. S2(d). After the peak of $S(\mathbf{q})$ shifts to the $\mathbf{\Gamma}$ point, its height increases within a specific range of $U/t$ [see, Fig. S4(b-d)], until reaching its maximum value, as illustrated by Fig. S3(c,d) and Fig. S4(c,d), indicating a transition from a partially-polarized to a fully-polarized ferromagnetic phase.
The above numerical observations reveal the presence of ferromagnetism with increasing Coulomb interaction. To confirm its robustness in the thermodynamic limit, we further employ the unrestricted Hartree-Fock mean-field approximation (see details in Supplementary Materials Section II). As shown in Fig. S2(g), the polarization $P\equiv 2Nm_{s}(\mathbf{\Gamma})$ increases gradually as $U/t$ becomes larger, and finally a full polarization state occurs at $U\simeq 4.5t$ for both $N=48\times 48$ and $N=24\times 24$ systems under periodic boundary conditions on both directions. While disparities in polarization $P$ between $N=48\times 48$ and $N=24\times 24$ lattice sizes are notable for $U\lesssim 4.5t$, beyond this threshold ($U\gtrsim 4.5t$) in the fully polarized ferromagnetic state, the polarization $P$ shows minimal sensitivity to the size of the system. This echos the observations in Fig. S2(e), suggesting the robust fully-polarized ferromagnetism regardless of system sizes in the strong coupling regime.

### v2

@@ Discussion
Fig. 6: $|$ Line-cut plot of the static spin structure $\boldsymbol{S(\mathbf{q})}$ with respect to various bond dimensions $\boldsymbol{D}$. The line-cut path is chosen along the momentum path passing through the $\Gamma$ point (denoted as the white dashed line in Fig. 1b in the main text). Parameters for system size NN, coupling $U/t$, electron doping $\delta$: (a) $N=24\times 6$, $U/t=10$, $\delta=1/2$; (b) $N=24\times 6$, $U/t=24$, $\delta=1/2$; (c) $N=8\times 4$, $U/t=6$, $\delta=1/2$; (d) $N=8\times 4$, $U/t=8$, $\delta=1/2$; (e) $N=24\times 4$, $U/t=\infty$, $\delta=3/16$; (f) $N=24\times 4$, $U/t=\infty$, $\delta=1/4$.
METHODS
Density matrix renormalization group
In the numerical part, we employ DMRG algorithm to simulate the Hubbard model Hamiltonian in Eq. (1) on triangular lattice cylinders with primitive vectors $\boldsymbol{e}_{x}=(1,0)$, $\boldsymbol{e}_{y}=(1/2,\sqrt{3}/2)$, wrapping on cylinders with a lattice spacing of unity. System size is $N=L_{x}\times L_{y}$, where $L_{x}$ ($L_{y}$) represents the length (circumference) of the cylinder. Here we mainly study cylinders with width $L_{y}=4$, $6$ and set the energy unit as $t$. We remark that, for the triangular lattice, $\mathbf{K}$ points are inaccessible on cylinders with width $L_{y}=4$. As a result, it is hard to exactly capture the magnetic orders with periodicity of 3, such as the $120^{\circ}$ Néel order. Therefore, we also examine the $L_{y}=6$ cylinders to suppress possible finite-size effects.
Our numerical simulation is based on the Hubbard model and our focus is the intermediate $U$ Hubbard model. Only in the infinite-$U$ limit, we use the effective $t$-$J$ model with vanishing superexchange interaction $J=4t^{2}/U$.
@@ METHODS
For DMRG calculations, given the varying convergence rates at different parameters, we set the DMRG bond dimension $D$ up to $D\approx 36,000$. Notably, large bond dimensions are necessary for distinguishing between ferromagnetic and incommensurate spin-density-wave (SDW) phases. We provide specific examples to illustrate the necessity of large bond dimensions for accurately identifying the ground-state properties of both the incommensurate SDW and ferromagnetic phases. Additionally, we have performed DMRG calculations with different initial states to ensure the convergence of DMRG. (1) For incommensurate SDW region, at relatively small bond dimensions, the peak of the static spin structure $S(\mathbf{q})$ is located at the momentum $\mathbf{\Gamma}$ [see Figs. 6a, c, d], which is a signature for ferromagnetism. However, with the increase of bond dimension, the energy becomes lower and the peak splits around the momentum $\mathbf{\Gamma}$. These observations indicate large bond dimensions are required to obtain converged results for distinguishing the ferromagnetism and iSDW with close energies. (2) On the other side, for the ferromagnetism region, a large bond dimension is crucial for determining whether the ground state is partially polarized or fully polarized. As presented in Fig. 6b, the peak of $S(\mathbf{q})$ gets higher with increasing bond dimension, until it saturates. Notably, DMRG convergence becomes more challenging with decreasing interaction strength $U/t$ and increasing system size $N$. In Figs. 6e, f, we also present the good convergence of the DMRG simulations in the iSDW region at $U=\infty$ that separates the ferromagnetic phases at $\delta\approx 1/16$ and $\delta\approx 1/2$ for $L_{y}=4$ [see Supplementary Note 4].
Unrestricted Hartree-Fock analysis
@@
In the analytical part, we apply the unrestricted Hartree-Fock (UHF) approximation. The mean-field order parameters are introduced for both the on-site densities $\langle n_{\mathbf{i},\sigma}\rangle$ and the spin-flips $\langle S_{\mathbf{i}}^{-}\rangle$ and $\langle S_{\mathbf{i}}^{+}\rangle$. We have a total $4N$ variational parameters with $N$ being the number of the lattice sites. These variational parameters can be iteratively solved by the self-consistent equations
$\langle n_{\mathbf{i},\sigma}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|n_{\mathbf{i},\sigma}|\psi_{n}\rangle,$ (11)
$\langle S_{\mathbf{i}}^{\pm}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|S_{\mathbf{i}}^{\pm}|\psi_{n}\rangle,$ (12)
@@
For a stable convergence, we utilize the direct inversion in the iterative subspace (DIIS) method. In this approach, we can express the Hartree-Fock Hamiltonian in the matrix form as follows:
@@
$H_{\mathrm{HF}}=C^{\dagger}FC+\sum_{\mathbf{i}}U\left[-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle\right],$ (13)
where $C$ is a vector with $C_{\mathbf{i}\sigma}=c_{\mathbf{i},\sigma}$ and $F$ is the Fock matrix. For a magnetic order, the single-particle correlation matrix $\Phi$ is defined as $\Phi_{\mathbf{ij},\sigma\sigma^{\prime}}=\langle c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma^{\prime}}\rangle.$ The DIIS method relies on the property that a convergence condition is met when the single-particle correlation matrix commutes with the Fock matrix $F\Phi-\Phi F=0.$ This condition allows us to introduce an error vector for the $n$-th iteration as $\mathbf{e}_{n}=F_{n}\Phi_{n}-\Phi_{n}F_{n},$ where $\Phi_{n}$ is obtained from the diagonalization of $F_{n}$. The iteration process is to update the Fock matrix $F_{n}$ at $n$th step as $F_{n}=\sum_{i=1}^{k}d_{n-i}F_{n-i},$ by involving the last $k$ Fock matrices $F_{n-i}$} ($i=1,\cdots,k$). Here, $d_{n-i}$ are the DIIS coefficients obtained through a least-squares constrained minimization of the error vectors. That is, we perform the optimization by minimizing the error $\epsilon=|\sum_{i=1}^{k}d_{n-i}\mathbf{e}_{n-i}|^{2}$ under the constraint $\sum_{i=1}^{k}d_{n-i}=1$. In our calculations, we set the convergent condition of the energy difference between two subsequent iteration processes as $10^{-8}t$.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.
CODE AVAILABILITY
All numerical codes in this paper are available upon request to the authors.
Acknowledgements
We thank helpful discussions with Ashvin Vishwanath and Yuchi He. This work is supported by the National Natural Science Foundation of China (Grant No.92477106) and the Fundamental Research Funds for the Central Universities.
Author contributions
All authors contributed to developing the ideas, analysing the results and writing the manuscript. Z.Z. initiated and led the project. Q.Q.C. performed the numerical calculations and data analysis of the main results. S.A.C. performed the mean-field analysis.
Competing interests
The authors declare no competing interests.
Supplementary Information for
“Geometric Frustration Assisted Kinetic Ferromagnetism in Doped Mott Insulators”
Qianqian Chen Shuai A. Chen Zheng Zhu
I Evolution of the magnetic structure factor at Electron Doping $\delta=1/2$
@@
Fig. S1: The contour plot of the static spin structure factor $\boldsymbol{S(\mathbf{q})}$. Calculated with $N=8\times 4$ at $\delta=1/2$ for the Hubbard model.
@@
Fig. S2: The contour plot of the static spin structure factor $\boldsymbol{S(\mathbf{q})}$. Calculated with $N=12\times 4$ at $\delta=1/2$ for the Hubbard model.
@@
We present additional supporting data for the evolution of the momentum $\mathbf{q}_{0}$ of the peak in $S(\mathbf{q})$ as a function of interaction strength $U/t$ in Fig. S1 for $N=8\times 4$ and Fig. S2 for $N=12\times 4$. These results for $L_{y}=4$ systems are consistent with the case of $L_{y}=6$ in Figs. 1(a)-(b) in the main text. The peak locations in $S(\mathbf{q})$ progressively shift toward the $\mathbf{\Gamma}$ point as $U/t$ increases, as presented in Fig. S1(a)-(c) and Figs. S2(a), (b). This observation supports the schematic plot in Fig. 1(c) in the main text. After the peak of $S(\mathbf{q})$ shifts to the $\mathbf{\Gamma}$ point, its height increases within a specific range of $U/t$ [see Figs. S2(b)-(d)], until reaching its maximum value, as illustrated by Figs. S1(c), (d), and Figs. S2(c), (d), indicating a transition from a partially-polarized to a fully-polarized ferromagnetic phase.
II Perturbation and the wave function analysis for ferromagnetic mechanism
As shown in the main text, the Hubbard Hamiltonian encompasses several types of hopping processes. Our measurements of each microscopic process reveal that the ferromagnetism at finite $U$ arises solely from the doublon-singlon exchange. Therefore, in the following, we concentrate on the doublon-singlon exchange, and aim to understand why ferromagnetism is present on the electron-doped cases but absent on the hole-doped cases. To address this, we have developed two lines of reasoning, i.e., the perturbation analysis and the wave function analysis.
Fig. S3: Schematic illustration of hopping processes in the electron-doped Hubbard model. Here, we have taken the denotation of the bases into consideration for the hopping signs.
For the perturbation analysis, we consider two types of effective spin exchange resulting from kinetic virtual processes: Superexchange, commonly encountered in the Hubbard model, giving rise to the term $\sim{t^{2}}/{U}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$; and doublon-assisted ring exchange, represented as $\sim t^{3}/\Delta E^{2}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, where $\Delta E$ is the energy difference between the initial and intermediate states. In the ring exchange process, a doubly occupied site is necessary for one electron to exchange while another passes by, as illustrated in Fig. 3(d) in the main text. Notably, $\Delta E$ is of the order of $t$ because the presence of two neighboring doublons restricts the hopping of singlons, preventing them from gaining kinetic energy. For electron doping (where $t<0$), the virtual processes involved in ring exchange $t^{3}/\Delta E^{2}$ promote ferromagnetism and dominates over the superexchange coupling $t^{2}/U$ for $U>t$, since from $\Delta E\sim|t|<{U}$ we have $|t^{3}/\Delta E^{2}|>t^{2}/U$. On the other side, for hole doping (where $t>0$), the ring exchange $t^{3}/\Delta E^{2}$ favor antiferromagnetism, similar to the effect of superexchange coupling.
Fig. S4: Schematic illustration of hopping processes in the hole-doped Hubbard model. Here, we have taken the denotation of the bases into consideration for the hopping signs.